### v1

Abstract
This paper presents a state-of-the-art algorithm for the vertex enumeration problem of arrangements, which is based on the proposed new pivot rule, called the Zero rule. The Zero rule possesses several desirable properties: i) It gets rid of the objective function; ii) Its terminal satisfies uniqueness; iii) We establish the if-and-only-if condition between the Zero rule and its valid reverse, which is not enjoyed by earlier rules; iv) Applying the Zero rule recursively definitely terminates in $d$ steps, where $d$ is the dimension of input variables. Because of so, given an arbitrary arrangement with $v$ vertices of $n$ hyperplanes in $\mathbb{R}^{d}$, the algorithm’s complexity is at most $\mathcal{O}(n^{2}d^{2}v)$ and can be as low as $\mathcal{O}(nd^{4}v)$ if it is a simple arrangement, while Moss’ algorithm takes $\mathcal{O}(nd^{2}v^{2})$, and Avis and Fukuda’s algorithm goes into a loop or skips vertices because the if-and-only-if condition between the rule they chose and its valid reverse is not fulfilled. Systematic and comprehensive experiments confirm that the Zero rule not only does not fail but also is the most efficient.
keywords:
Hyperplane arrangement; vertex enumeration; polynomial pivot rule.
1 Introduction
Vertex enumeration (VE) is a fundamental problem identifying vertices that satisfy specific linear constraints [5, 9, 15, 16, 26, 35]. When these constraints are linear inequalities like $Ax\leq b$, it entails enumerating vertices on a polytope, while constraints such as $Ax=b$ correspond to enumerating vertices on an arrangement. Geometrically, VE bridges two crucial representations of a polytope or an arrangement: $H$-representation and $V$-representation, which respectively define a polytope or an arrangement from its faces and vertices. As such, VE is not only a fundamental problem in computational discrete geometry but also holds a wide array of real-world applications. VE has been applied in problems including, among others, effective field theory [39], implicit mesh generation [23], and computing approximating polytopes [3]. More recently, VE has emerged as a key component in unraveling the inner workings of deep learning [17, 37, 38]. The underlying concept stems from the fact that a network with piecewise linear activation functions serves as a piecewise linear entity, partitioning the space into numerous linear regions. Each region is defined by a set of linear constraints. Identifying the vertices of these linear constraints is pivotal in understanding the dynamics and robustness of deep networks [38].
The earliest published method to solve the problem of VE can be found in [26]. Their approach is to add inequalities in a stepwise manner. In each step, “new” extreme points are created and “old” ones are excluded. Subsequently, due to the intrinsic connection between VE and linear programming, the simplex method and its variants, such as Abdullahi and Chen’s algorithm [1], Avis and Fukuda’s pivoting method [4], Balinski’s algorithm [6], Bremner’s primal simplex method [9], Chand and Misra’s algorithm [10], Chvatal’s linear programming approach [11], Dyer and Frieze’s complexity analysis [15], Elbassioni and Raychaudhuri’s enumeration algorithm [16], Mattheiss’s approach [24], Provan’s efficient algorithm [28], Shioura’s optimal algorithm [31], were utilized to solve the VE problem. The simplex method was originally used in linear programming, based on the observation that a linear objective function always attains its maximum at extreme points of the feasible region (polytope). Geometrically, one only needs to move the current vertex to an adjacent one, and if the maximum exists, the algorithm will terminate at one vertex that maximizes the objective function. Specifically, in this method, a dictionary is used to represent the current vertex, and a step called “pivoting” is employed to select entering and leaving variables to update the dictionary, thus traversing between neighboring vertices. However, if a vertex lies on more than $d$ hyperplanes, the pivoting may trap into several dictionaries concerning the same vertex, which greatly deteriorates the efficiency. Hence, different pivot rules were introduced to avoid this situation and boost efficiency. For example, the Bland’s rule [7, 8], the Edmonds-Fukuda rule [18] and its variant [12], the Criss-Cross rule [32, 33, 36] and its variant [19], and the Jensen’s general relaxed recursion [21].
Many previous papers were dedicated to the VE of a polytope. Although the importance of VE of an arrangement is increasingly recognized, to the best of our knowledge, only a few methods focus on VE of an arrangement. The method in Moss [25] is designed for listing exactly one co-basis from each orbit under the action from a set of automorphisms to the arrangement. It considers the smallest positive value and the largest negative value in the ratio test of every column, thereby obtaining all adjacent dictionaries. By repeating this process, every dictionary is obtained. Then, the algorithm compares each dictionary in this arrangement to determine whether it is a duplicate. This method is essentially a brute-force enumeration, with a high demand for storage and computational complexity. The method proposed by Avis and Fukuda [4] is based on an observation that since repeating one pivot rule can yield a path from some dictionary to an optimal dictionary, then tracking back the path can generate all dictionaries ending at that optimal dictionary, without the need for the brute-force enumeration. Instead of directly computing coordinates of all vertices, the algorithm outputs lexicographically minimum basis for each vertex, and then coordinates can be computed based on the outputs if needed.
Despite the attraction of the idea, the algorithm in Avis and Fukuda [4] should be upgraded from the following three major aspects: i) The chosen Criss-Cross rule and its valid reverse, unfortunately, do not satisfy a necessary and sufficient (if-and-only-if) condition, which fails the searching because it makes the pivoting either go to a loop or catastrophically skip some vertices. ii) The objective function is not necessary for the VE of an arrangement. Firstly, the objective function restricts the traversal of vertices, leading to more steps in complex arrangements. Secondly, the value of the objective function is unpredictable until reaching a vertex, requiring considering all vertices to optimize the objective function, which is more time-consuming. Thirdly, the pivot operation on a dictionary can be performed without the objective function, and neither the magnitude of the objective function nor the objective function itself affects our computation of the vertex coordinates based on the dictionary’s basis or co-basis. Therefore, the objective function only serves as a guide during the pivot process, but in our study, this guidance can also be observed from the dictionary itself. iii) The Criss-Cross rule may not have a unique convergence, which introduces more operations in the VE problem and makes the algorithm difficult to organize.
In this paper, building upon the framework proposed in Avis and Fukuda [4], we develop a new pivot rule that is specifically tailored for the VE of an arrangement. This rule is not a small modification but solves the key issues of the Criss-Cross rule and leads to a more efficient algorithm. The new pivot rule, dubbed as the Zero rule, uses the linear independence of normal vectors among hyperplanes in the arrangement. The most distinct features of the Zero rule in contrast to the Criss-Cross rule are that i) the if-and-only-if condition between the Zero rule and its valid reverse is established, which can ensure all vertices are discovered; ii) it gets rid of objective functions; iii) its convergence is unique, and the number of convergence steps is bounded by the dimension $d$, which arranges all vertices in a single and shallow tree, thereby saving computation. Given a simple arrangement with $v$ vertices of $n$ hyperplanes in $\mathbb{R}^{d}$, the complexity of the algorithm using the Zero rule is at most $\mathcal{O}(n^{2}d^{2}v)$ but can be as low as $\mathcal{O}(nd^{4}v)$ for simple arrangements. Table 1 summarizes the characteristics of Moss [25], Avis and Fukuda [4], and our proposed algorithm. Furthermore, we provide systematic and comprehensive experiments to compare the performance of different algorithms. Experiments confirm that our proposed algorithm using the Zero rule not only does not fail but also is the most efficient.
Table 1: A summary between Moss [25], Avis and Fukuda [4] and ours, where $n$ is number of hyperplanes, $d$ is the dimension, and $v$ is the number of vertices.
storage obj. func. no cycle complexity single tree
[25] $\mathcal{O}(ndv)$ $\mathcal{O}(nd^{2}v^{2})$
[4] $\mathcal{O}(nd)$
Ours $\mathcal{O}(1)$ $\mathcal{O}(nd^{4}v)$ or $\mathcal{O}(n^{2}d^{2}v)$
2 Notations and Definitions
For convenience and consistency, notations in this paper are the same as [4] that uses dictionaries instead of the simplex tableau.
Definition 2.1 (Hyperplane and arrangement).
[…]
Definition 2.2 (Vertex of an arrangement).
A vertex of an arrangement is the unique solution to the system of $d$ equations corresponding to $d$ intersecting hyperplanes. Moreover, a vertex is degenerate if it is contained in more than $d$ hyperplanes.
The vertex enumeration problem for a hyperplane arrangement is enumerating all vertices of the arrangement.
Definition 2.3 (Dictionary with the objective function).
Suppose there are $n$ hyperplanes in an arrangement, denoted by slack variables $x_{i}=b_{i}x_{g}-\langle\mathbf{c}_{i},\mathbf{y}\rangle,i=1,\cdots,n$, where $x_{g}$ has a constant value $1$. Let $x_{f}$ be the objective function which is the linear combination of elements in $\mathbf{y}$. For a set of linearly independent vectors $\{\mathbf{c}_{j_{1}},\cdots,\mathbf{c}_{j_{d}}\}$, $y_{1},\cdots,y_{d}$ can be expressed in terms of $\mathbf{x}_{j_{1}},\cdots,\mathbf{x}_{j_{d}}$, leading to a relation between slack variables:
$\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N},$ (1)
where $\mathbf{x}_{B}=(x_{i_{1}},\cdots,x_{i_{n-d}},x_{f})$ and $\mathbf{x}_{N}=(x_{g},x_{j_{1}},\cdots,x_{j_{d}})$. Such a relation is called a dictionary. A dictionary is degenerate if it is associated with a degenerate vertex. Figure 1 shows a dictionary with the objective function.
Definition 2.4 (Basis and co-basis).
In a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, the set $B=\{i_{1},\cdots,i_{n-d},f\}$ is called basis, and $\mathbf{x}_{B}$ are basic variables. $N=\{g,j_{1},j_{2},\cdots,j_{d}\}$ is called co-basis, and $\mathbf{x}_{N}$ are non-basic variables. $\bar{A}=(\bar{a}_{ij})\in\mathbb{R}^{{(n-d+1)}\times{(d+1)}}$ is the coefficient matrix, and $\bar{a}_{ij}$ expresses entries in $\bar{A}$ corresponding to $x_{i}$ and $x_{j}$.
Definition 2.5 (Dictionary without the objective function).
As Figure 1 shows, a dictionary can go without the objective function. Accordingly, simply removing the last row of Eq. (1), $\mathbf{x}_{B}=(x_{i_{1}},\cdots,x_{i_{n-d}})$ and $B=\{i_{1},\cdots,i_{n-d}\}$.
In the proposed Zero rule, all dictionaries work without the objective function. Each dictionary is associated with one vertex, while each non-degenerate vertex is associated with only one dictionary. The vertex enumeration problem can be transformed into enumerating all dictionaries. The key in this process is to circumvent yielding vertices associated with multiple dictionaries more than once.
Definition 2.6 (Primal & Dual Feasible [4]).
In a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, if $\bar{a}_{ig}\geq 0$, then the variable $x_{i},i\in B_{\neq f}$ is primal feasible. If $\bar{a}_{fj}\leq 0$, then the variable $x_{j},j\in N_{\neq g}$ is dual feasible. A dictionary is primal feasible, if $\forall i\in B_{\neq f},x_{i}$ is primal feasible. A dictionary is dual feasible, if $\forall j\in N_{\neq g},x_{j}$ is dual feasible. The dictionary is optimal, if it is both primal and dual feasible. Note that these definitions are only applicable to the dictionary with the objective function.
Figure 1: A dictionary with and without the objective function.
Definition 2.7 (Pivoting).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary, a pivot $(r,s)$ on the dictionary is an exchange between $x_{r}$ in $\mathbf{x}_{B}$ and $x_{s}$ in $\mathbf{x}_{N}$, which will generate a new dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, where the new coefficient matrix $\tilde{A}=(\tilde{a}_{ij})$ is obtained by computing Eq. (2) and sorting rows and columns of $\tilde{A}=(\tilde{a}_{ij})$ in increasing order of indices:
$\tilde{a}_{sr}=\cfrac{1}{\bar{a}_{rs}},$ $\tilde{a}_{ir}=\cfrac{\bar{a}_{is}}{\bar{a}_{rs}},$ $\tilde{a}_{sj}=-\cfrac{\bar{a}_{rj}}{\bar{a}_{rs}},$ $\tilde{a}_{ij}=\bar{a}_{ij}-\cfrac{\bar{a}_{is}\bar{a}_{rj}}{\bar{a}_{sr}},$ $i\in B_{\neq r},j\in N_{\neq s}.$ (2)
Definition 2.8 (Valid reverse pivoting).
Let $D$ be an arbitrary dictionary, a reverse pivot $(s,r)$ on $D$ will return a new dictionary $\tilde{D}$. It is a valid reverse about the rule $\mathcal{R}$ if applying $\mathcal{R}$ on $\tilde{D}$ fixed at $(s,r)$ instead of other entries.
Definition 2.9 (Lexicographically minimum basis [4]).
Let $B_{1},B_{2},\cdots,B_{n}$ be basis for all dictionaries associated with the same vertex. The basis attaining the minimum lexicographic order among $B_{1},B_{2},\cdots,B_{n}$ is called the lexicographically minimum basis. Verifying the lexicographically minimum basis and only computing the vertex using this dictionary can avoid producing the same vertex more than once.
It is important to note that for a non-degenerate vertex associated with only one dictionary, this dictionary (non-degenerate dictionary) is always lexicographically minimum. Testing if the basis of a degenerate dictionary is lexicographically minimum is called the lex-min test. Let $B$ be a basis for a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$. $B$ is not lexicographically minimum if and only if there exists $r\in B_{\neq f}$ and $r\in N_{\neq g}$ such that $r>s,\bar{a}_{rg}=0$ and $\bar{a}_{rs}\neq 0$ [4].
The lex-min test can avoid outputting a vertex more than once: First, given a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, let $B_{1}=\{i\in B_{\neq f}~{}|~{}\bar{a}_{ig}=0\}$, which selects the indices whose associated constant columns are zero. Then, $x_{i}=0$ for $i\in B_{1}$, which means that all the $i$-th hyperplanes, $i\in B_{1}\cup N$, pass through the same vertex. Consequently, the $i$-th hyperplanes, $i\notin B_{1},i\in B$ will never pass through this vertex. Thus, checking rows covered by $B_{1}$ is sufficient to analyze, without the need for really solving the dictionary. If a dictionary $D$ is not lex-min (e.g. LHS of Figure 2), then there will always exist another dictionary $\tilde{D}$ in the entire arrangement (e.g. RHS of Figure 2), where the lexicographical order of $\tilde{B}$ is smaller than $B$. Since the lexicographically smallest dictionary is unique at each vertex, we can ensure that each vertex is counted only once, thereby saving the computational overhead.
Figure 2: An example of the lex-min test. LHS: not pass the test. RHS: pass the test.
Definition 2.10 (The Criss-Cross rule).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary with an objective function. The Criss-Cross rule [32, 33, 36] is as follows: 1) $i\neq f,g$ is the smallest index such that $x_{i}$ is primal or dual infeasible. 2) If $i\in B_{\neq f}$, let $r=i$ and $s$ be the minimum index such that $\bar{a}_{rs}>0$; otherwise, let $s=i$ and let $r$ be the minimum index such that $\bar{a}_{rs}<0$. 3) Pivot $(r,s)$ and update the dictionary.
The condition to verify valid reversibility of the Criss-Cross rule that was proposed in [4] is that if $(s,r)$, where $s\in B_{\neq f}$ and $r\in N_{\neq g}$, is a valid reverse of the Criss-Cross rule for a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, then either (a) $\bar{a}_{sg}>0,\bar{a}_{sr}>0,\bar{a}_{sj}$ for $j\in N_{\neq g},j<s$, or (b) $\bar{a}_{fr}<0,\bar{a}_{sr}<0,\bar{a}_{ir}<0~{}\mathrm{for}~{}i\in B_{\neq f},i<r$.
Theorem 2.11.
The condition provided by Avis and Fukuda in [4] is a necessary but not sufficient condition of valid reversibility of the Criss-Cross rule.
Proof: Assuming that from the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, by pivoting $(s,r)$, we obtain $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$. The condition (a) ensures the selection of $s$ in the second step of the Criss-Cross rule, which gives a specific row $r\in\tilde{B}$. But it cannot determine whether the first step of the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ identifies the desired $r$-th row. Similarly, the condition (b) guarantees the selection of $r$ in the second step of the Criss-Cross rule, which is given by a specific row $s\in\tilde{N}$. However, it cannot guarantee whether the first step of the Criss-Cross rule will yield the required column $r$. Therefore, even if $(s,r)$ in $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ satisfies their given conditions, it still cannot guarantee that applying the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ will pinpoint $(r,s)$.
$\Box$
3 The Zero Rule
In this section, we propose the Zero rule which works on the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ without the objective function:
1) Let $s$ be the smallest index in $N_{\neq g}$ such that $\{i\in B~{}|~{}\bar{a}_{is}\neq 0\textup{ and }i<s\}\neq\emptyset$.
2) After $s$ is selected, let $r=\min\{i\in B~{}|~{}\bar{a}_{is}\neq 0\textup{ and }i<s\}$.
3) Pivot $(r,s)$ and update the dictionary.
The Zero rule represents an advance beyond the Criss-Cross rule, no longer focusing on the objective function but on the entries of the coefficient matrix, which is the arrangement itself. Now, we prove the desirable properties of the proposed Zero rule. First, we show that given a dictionary, the Zero rule will only select a unique entry to pivot, which is obtained by direct calculus.
Proposition 1 (Uniqueness).
The Zero rule only selects a unique entry of a dictionary to pivot.
Proof: Direct calculus.
$\Box$
Next, we give a condition to test if an entry of a dictionary passes the valid reverse pivoting test. The criterion is that if $(s,r)$ passes the test, then applying the Zero rule in the new dictionary will automatically pinpoint $(r,s)$. Then, pivoting $(r,s)$ will go back to the original dictionary. This property is called reversibility, which is important to the vertex enumeration algorithm in terms of ensuring that no storage is needed, and no vertices are missed. Figure 3 presents an example to illustrate the Zero rule and its valid reverse, where blocks of different colors correspond to different sub-conditions taking effect.
Proposition 2 (Reversibility).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be an arbitrary dictionary. Then $(s,r)$, where $s\in B$ and $r\in N_{\neq g}$, is the valid reverse of the Zero rule if and only if
i) $r<s$ and $\bar{a}_{sr}\neq 0$;
ii) $\forall j\in N_{\neq g}$ and $r<j<s$, $\bar{a}_{sj}=0$;
iii) $\forall j\in N_{\neq g}$ and $j<s$, for those $i\in B$ and $i<j$, $\bar{a}_{ij}=0$.
Figure 3: An exemplary illustration of the Zero rule and its valid reverse. Blocks of different colors represent where different sub-conditions are enforced.
Proof: Let $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ be the dictionary obtained from $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ through pivoting $(s,r)$. Now, let us prove the if-and-only-if condition, respectively.
$\Rightarrow$: Assume $(s,r)$ is the valid reverse pivot of the proposed Zero rule, the resultant dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ will make the proposed Zero rule automatically pinpoint $(r,s)$ to pivot. This means that 1) $\forall j\in\tilde{N}_{\neq g}$ and $j<s$, any $i\in\tilde{B}_{\neq r}$ and $i<j$, $\tilde{a}_{ij}=0$; 2) $\tilde{a}_{rj}=0$ for any $r<j<s$; otherwise $(r,j)$ will be selected to pivot. 3) $\tilde{a}_{rs}\neq 0$ and $r<s$. Now, let us use these three properties to deduce conditions i)-iii).
$\bullet$ The condition i) holds from 3) directly.
$\bullet$ As far as the condition ii) is concerned, according to 2), we have $\tilde{a}_{rj}=0$ for any $r<j<s$. Combining $\tilde{a}_{rj}=\cfrac{\bar{a}_{sj}}{\bar{a}_{sr}}=0$ leads to that $\bar{a}_{sj}=0$ for $r<j<s$.
$\bullet$ 1) is almost the condition iii). Combining Eq. (2) and ii), $\tilde{a}_{ij}=0$ can naturally lead to $\bar{a}_{ij}=0$.
$\Leftarrow$: Assume conditions i)-iii) hold, we have
$\bullet$ The condition i) ensures the entry in $(r,s)$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ is non-zero and $r<s$.
$\bullet$ As far as 3) is concerned, according to the condition ii), we have $\bar{a}_{sj}=0$ for any $r<j<s$. Combining $\tilde{a}_{rj}=\cfrac{\bar{a}_{sj}}{\bar{a}_{sr}}=0$ leads to that $\tilde{a}_{sj}=0$ for $r<j<s$.
$\bullet$ The condition iii) is almost 1). By Eq. (2) and ii), $\bar{a}_{ij}=0$ can naturally lead to $\tilde{a}_{ij}=0$.
$\Box$
After deriving the valid reverse of the Zero rule, we introduce another feature of indices leaving the co-basis when repetitively enforcing the Zero rule.
Lemma 3.1 (Index increasing).
Starting with any dictionary arrangement, when the Zero rule is repeated, the sequence of indices leaving the co-basis in each step increases.
Proof: Assume in the $l$-th step, the Zero rule is fixed at $(r_{l},s_{l})$ and yields the $l$-th dictionary $\mathbf{x}_{B_{l}}=\bar{A}\mathbf{x}_{N_{l}}$, where the Zero rule pinpoints $(r_{l+1},s_{l+1})$.
$Zero\ rule]{(r_{l},s_{l})}\mathbf{x}_{B_{l}}=\bar{A}_{l}\mathbf{x}_{N_{l}}\xrightarrow[Zero\ rule]{(r_{l+1},s_{l+1})}\cdots$
Note that $(s_{l},r_{l})$ is the valid reverse pivot of the proposed Zero rule, then by Proposition 2, we have $\forall j\in N_{\neq g}\textup{ and }j<s_{l},\textup{ for those }i\in B\textup{ and }i<j,\bar{a}_{ij}=0$.
For columns in $\mathbf{x}_{B_{l}}=\bar{A}\mathbf{x}_{N_{l}}$, from ii, iii), $\forall j\in N_{\neq g}\textup{ and }j<s_{l}$, $\{i\in B_{l}~{}|~{}i<j,\bar{a}_{ij}\neq 0\}=\emptyset$. At the same time, $x_{s_{l}}$ has been moved into $\mathbf{x}_{B}$. Thus, applying the Zero rule again would cause the later column $s_{l+1}>s_{l}$ to be selected. Hence, the index leaving the co-basis is increasing.
$\Box$
Based on Lemma 3.1, one can prove that the number of pivot steps of the Zero rule is no more than $d$.
Proposition 3 (Termination in $d$ steps).
For any dictionary in an arrangement of $\mathbb{R}^{d}$, consecutively applying the Zero rule or its reverse will attain the terminal in at most $d$ steps.
Proof: According to Lemma 3.1, it can be concluded that each time the Zero rule is applied, the column it selects will be at least one position later compared to the previous step. In any dictionary in $\mathbb{R}^{d}$, there are a total of $d$ columns, excluding the column corresponding to $x_{g}$. This implies that the Zero rule will be repeated for at most of $d$ times.
$\Box$
The property given in Proposition 3 is very important because most pivot rules, such as Bland’s rule and the Criss-Cross rule, converge in exponential steps. The exponential convergence of Bland’s rule is proven in [27, 34], while the Criss-Cross rule is shown by an example that requires at least $12^{n}-$ 1 steps for convergence [29]. Moreover, a better method, such as the Random-Facet pivot rule mentioned in [14, 20], takes at most a subexponential number of pivot steps. Therefore, the convergence speed of our Zero rule is very fast.
After studying the number of convergence steps used in the proposed Zero rule, we focus on the number of terminations among the arrangement. Will the pivot rule lead to multiple terminations? To address this question, let us first identify what “0” means in each dictionary.
Lemma 3.2.
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be an arbitrary dictionary. Then for any $i\in B$, the normal vector of the $i$-th hyperplane $\mathbf{c}_{i}\in\mathbb{R}^{d}$ satisfy $\mathbf{c}_{i}\in\mathrm{span}(\{\mathbf{c}_{j}~{}|~{}\bar{a}_{ij}\neq 0\})$.
Proof: In any fixed dictionary $x_{B}=\bar{A}x_{N}$, $x_{i}=\sum\limits_{j\in N}\bar{a}_{ij}x_{j}$. Also, note that we have
$\begin{cases}&x_{i}=b_{i}x_{g}-\langle\mathbf{c}_{i},\mathbf{y}\rangle,\\ &x_{j}=b_{j}x_{g}-\langle\mathbf{c}_{j},\mathbf{y}\rangle,\end{cases}$ (3)
for all $i\in B$ and $j\in N,j\neq g$. Substituting Eq. (3) into $x_{i}=\sum\limits_{j\in N}\bar{a}_{ij}x_{j}$, we can derive the relationship between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ as follows:
$c_{il}=\sum\limits_{j\in N_{\neq g}}\bar{a}_{ij}c_{jl},\qquad\qquad l\in\{1,2,...,d\},$ (4)
where $c_{il}$ and $c_{jl}$ are the $l$-th coordinate of the vector $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$, respectively. Wrapping it into a vector representation, we have
$\mathbf{c}_{i}=\sum\limits_{j\in N_{\neq g}}(\bar{a}_{ij}\mathbf{c}_{j})=\sum\limits_{j\in N_{\neq g},\bar{a}_{ij}\neq 0\ }(\bar{a}_{ij}\mathbf{c}_{j}),$ (5)
which means $\mathbf{c}_{i}\in\mathrm{span}(\{\mathbf{c}_{j}~{}|~{}\bar{a}_{ij}\neq 0\})$.
$\Box$
From the above lemma, it can be seen that the proposed Zero rule is not an indexing game. Instead, it is deeply related to the properties of the arrangement itself, as represented by the dictionary. The relationship between zero and non-zero entries is the reason why we call our proposed rule the Zero rule. Now, we show that under the Zero rule, there only exists one termination.
Proposition 4 (Unique convergence).
Repeating the Zero rule from any dictionary in the arrangement will result in a unique termination.
Proof: Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be the dictionary such that $N$ has attained the smallest lexicographical order among all co-basis in the system; note that the dictionary is a terminal of the Zero rule. Furthermore, suppose there exists another terminal of the Zero rule dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ and $\tilde{N}\neq N$, that is,
$\forall s\in\tilde{N}_{\neq g},\forall i<s\textup{ and }i\in\tilde{B},\tilde{a}_{is}=0.$ (6)
Comparing $N$ and $\tilde{N}$ entry by entry, let the $j$-th pair be the first such that they are not equal, and denote them as $n_{j}\in N,\tilde{n}_{j}\in\tilde{N}$.
Then, it is possible to construct a partition on $N$ and $\tilde{N}$, given as $N=N_{0}\cup\{n_{j}\}\cup N_{1},\tilde{N}=\tilde{N}_{0}\cup\{\tilde{n}_{j}\}\cup\tilde{N}_{1}$ where $N_{0},\tilde{N}_{0}$ are the set of all numbers before $n_{j}$ and $\tilde{n}_{j}$ in $N$ and $\tilde{N}$ respectively. By the definition of $n_{j}$ and $\tilde{n}_{j}$, $N_{0}=\tilde{N}_{0}$. $N_{1},\tilde{N}_{1}$ are remainders in $N$ and $\tilde{N}$, respectively.
Note that $\tilde{N}\neq N$, and $N$ attains the lexicographically minimum among all co-basis in the system, then $n_{j}<\tilde{n}_{j}$, and $\forall\tilde{n}_{l}\in\tilde{N}_{1},n_{j}<\tilde{n}_{j}<\tilde{n}_{l}$. In the dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ by Eq. (6), $\bar{a}_{n_{j}\tilde{n}_{j}}=0$ and for all $\tilde{n}_{l}\in\tilde{N}_{1}$, $\tilde{a}_{n_{j}\tilde{n}_{l}}=0$, which is shown as the Figure 4.
Figure 4: Zeros in the row of $x_{n_{j}}$ on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$.
It follows that $\{\mathbf{c}_{\tilde{n}_{l}}~{}|~{}\tilde{a}_{n_{j}\tilde{n}_{l}}\neq 0\}\subseteq\{\mathbf{c}_{\tilde{n}_{l}}~{}|~{}\tilde{n}_{l}\in N_{0}\}$. By Eq. (5), we have
$\mathbf{c}_{n_{j}}\in\mathrm{span}(\{\mathbf{c}_{\tilde{n}_{l}}~{}|~{}\bar{a}_{n_{j}\tilde{n}_{l}}\neq 0\})\subseteq\mathrm{span}(\{\mathbf{c}_{l}~{}|~{}l\in N_{0}\}).$ (7)
Since $N$ is a co-basis in the arrangement, then $\{\mathbf{c}_{l}~{}|~{}l\in N_{0}\}\cup\{\mathbf{c}_{n_{j}}\}\subseteq\{\mathbf{c}_{l}\ ~{}|~{}\ l\in N\}$ is linearly independent, which contradicts with Eq. (7). Hence, in the first place, any dictionary with a co-basis not attaining the lexicographical minimum order is not the end of pivoting by the Zero rule. It implies that the termination of repeating this rule from all dictionaries among an arrangement is unique, which is the co-basis with the lexicographical minimum order in the arrangement.
$\Box$
Connection to an open question. Propositions 3 and 4 might be related to an open question: “Is it possible to find a polynomial pivot rule for linear programming, i.e., a pivot rule which bounds the number of pivot steps by a polynomial function of the number of variables and the number of inequalities, or to prove that such a pivot rule does not exist”. This question is important [34], motivated by analyzing the complexity of the simplex method in linear programming. The diameter of a polytope provides a lower bound on the number of steps needed by the simplex method, although it is actually weaker than the so-called Hirsch conjecture in polyhedral theory. The Hirsch conjecture, formulated by W. M. Hirsch in 1957 in [22] and reported in the 1963 book of Dantzig [13], states that the edge-vertex graph of an $n$-facet polytope in $d$-dimensional Euclidean space has a diameter no more than $n-d$. This means that any two vertices of a polytope must be connected by a path of length at most $n-d$. However, this conjecture has been refuted by a counterexample presented in [30] which introduces a 43-dimensional polytope of 86 facets with a diameter exceeding 43. Additionally, a formal disproof is also outlined in [2].
However, this counterexample has no direct consequences for the analysis of the simplex method, as they do not rule out the possibility of a larger but still polynomial number of steps. While the proposed Zero rule is a polynomial pivot rule and is bounded by a linear function, it is on an arrangement instead of a polytope. Nevertheless, the Zero rule can shed some light on the open question. The Zero rule is based on the dictionary itself and goes without a ratio test. Thus, it does not inherently traverse between adjacent vertices or reflect the number of edges between vertices. If a ratio test could be incorporated or primal feasibility could be maintained within the Zero rule, it might serve as a polynomial pivoting rule on polytopes.
4 VE Algorithm Using the Zero Rule
In the above section, we introduce the Zero rule and its favorable properties. Here, we design the vertex enumeration algorithm with the Zero rule as its core technique. The designed algorithm inherits the beautiful framework in Avis and Fukuda [4] including the lex-min test and applying a rule and its valid reverse, among others. The core is we replace the Criss-Cross rule therein with the proposed Zero rule. Below we describe the proposed algorithm, elucidate its advantages, and conduct a detailed complexity analysis.
4.1 Algorithm Description
The pivoting method involves two main steps: initialization and search [4]. The initialization finds a vertex from a given arrangement and constructs an initial dictionary based on this vertex. Specifically, we take $d$ hyperplanes that intersect at a vertex. Then, we arrange the associated slack variables as co-basis variables and index them from $1$ to $d$. The rest slack variables will be basis variables and indexed from $d+1$ to $n$. According to Proposition 4, the so-initialized dictionary is the termination dictionary of all other dictionaries, as it is the lexicographically minimum.
Second, searching all dictionaries requires testing valid reverse for every entry of a dictionary until the if-and-only-if condition is satisfied. The search is based on the order of increasing index. When an entry fulfills the if-and-only-if condition, we pivot it and update the dictionary. Recursively conducting the reverse until the pivot rule is not fixed at any entry, which means that the algorithm arrives at the termination dictionary. At this time, as Figure 5 shows, we apply the Zero rule to go back to the previous dictionary and search the entry again. This will not cause repetition because the if-and-only-if condition can inform us of the previously visited entry. We can search for entries with higher indexes starting from the previously visited one.
Moreover, in the searching step, for each obtained dictionary, the lex-min test is done to verify if the dictionary has a lexicographically minimum basis. Only the basis of the dictionary that passes the test will output the vertex. This operation is to avoid producing the same vertex more than once.
By applying the Zero rule and its valid reverse, the searching path is a single spanning tree, and any two branches do not intersect, as Figure 5 shows. Each node in the spanning tree is guaranteed to be unique. However, Avis and Fukuda [4] did not provide the if-and-only-if condition, which may cause loops in the searching path. Moreover, by Proposition 4, there is only one dictionary that can not be applied to the Zero rule, and we do not need to store any intermediate dictionary. A detailed algorithm is presented as Algorithm $1$ in Figure 6.
Figure 5: Explination of the VE algorithm.
Since the Zero rule can converge in $d$ steps, one can set the initial dictionary to the 00-th layer. Each reverse pivot increments the layer by one, while each application of the Zero rule decrements the layer by one. When studying the dictionary of the $(d-1)$-th layer, we only need to determine whether the corresponding dictionary of the $d$-th layer satisfies the lex-min condition instead of examining the valid reverse. Then one can track back to the $(d-1)$-th layer. This can further save time. Therefore, our algorithm needs to save the layer number which is of $O(1)$ complexity.
Figure 6: The VE algorithm of an arrangement using Zero rule.
4.2 Advantage Analysis
Since our focus is VE of an arrangement, we only compare with the Criss-Cross rule and Jensen’s general relaxed recursion that are applicable to an arrangement. The advantages of the Zero rule are as follows:
Jensen’s general relaxed recursion requires partitioning a dictionary and analyzing it case by case. Its termination dictionary could be primal feasible, dual feasible, or optimal [21, 34], which makes it tedious for VE. As for the Criss-Cross rule, although it also needs to discuss two cases, it has only one type of termination dictionary, namely the optimal dictionary. This makes it better than Jensen’s general relaxed recursion, and it is also exactly the pivot rule used by Avis and Fukuda [4].
However, since the condition by Avis and Fukuda is not if-and-only-if, which means that after performing reverse pivoting, applying the Criss-Cross rule may not return to the original dictionary directly. As Figure 7 shows, when the reverse-applying operation about the Criss-Cross rule goes to a new entry, if this new entry has a smaller index, the algorithm goes into a loop. If this new entry has a larger index, the algorithm catastrophically skips some vertices. Both cases make the vertex enumeration fail. The advantage of using the if-and-only-if condition when testing valid reverse pivoting in vertex enumeration is that it guarantees one can return to the original dictionary without requiring additional storage. Since it is automatically known that the $(s,r)$ entry in the original dictionary has been pivoted, one can search for entries greater than $s$ and $r$ in further study.
Figure 7: The Criss-Cross rule may undergo loops and catastrophic skipping.
Moreover, compared to the Zero rule, the Criss-Cross rule is not ideal in both initialization and valid reverse. During the initialization process, the Zero rule is more convenient in generating all dictionaries where the pivot rule cannot be applied. By Proposition 4, through our way of initialization, the resultant dictionary with co-basis $\{g,1,2,\cdots,d\}$ is always the unique terminal dictionary. On the other hand, using Criss-Cross rules requires setting an objective function and generating all optimal dictionaries, which adds an extra step to the initialization process.
When it comes to the valid reverse, the advantages of the Zero rule are more obvious. The if-and-only-if condition between the Zero rule and its valid reverse is established so that all vertices are surely enumerated. But the Criss-Cross rule has a risk of being failed such as going into loops or skipping vertices. On the other hand, according to Proposition 3, the number of pivot steps of the Zero rule is bounded by $d$, whereas the pivot step of the Criss-Cross rule is exponential [29, 34]. Using the Zero rule, all dictionaries can be arrived in at most $d$-step reversals. At the same time, since the Zero rule can reach the unique termination in $d$ steps, using the Zero rule can skip testing valid reverse on dictionaries at the $d$-th layer, whereas the Criss-Cross rule requires studying every dictionary.
4.3 Complexity Analysis
For a fixed dimension $d$, we assume that there are $n$ hyperplanes, $v$ vertices in the arrangement, and $v_{d}$ vertices whose associated dictionary needs to apply the Zero rule $d$ times to arrive. When calculating a new dictionary, one needs $(n-d)(d+1)=\mathcal{O}(nd)$ calculations. Testing the lex-min also needs $\mathcal{O}(nd)$ calculations.
The computational complexity of the VE algorithm using the Criss-Cross rule ideally is only $\mathcal{O}(n^{2}dv)$, which has been provided in [4]. However, the rule used in [4] does not fulfill the if-and-only-if condition, which means that it may trap into loops when encountering degenerate vertices or skip certain vertices. We have constructed an example to illustrate this point later. Thus, the actual computational complexity is far worse than $\mathcal{O}(n^{2}dv)$ in some cases.
Now, let us analyze the complexity of the method from Moss [25]. In each dictionary, one needs to compute the ratio of the column of $x_{g}$ with every column and then select the extreme value. This step incurs a cost of $\mathcal{O}(d(n-d)v)=\mathcal{O}(ndv)$. Subsequently, during the comparison with discovered dictionaries, their co-basis need to be checked. If there are $k$ recorded dictionaries, then comparison requires $\mathcal{O}(kd)$ for each candidate. During the procedure of generating the adjacent co-basis about a fixed dictionary, one may find at most $nd$ candidates from it, comparing each candidate takes $\mathcal{O}(knd^{2})$ computations. Since there are a total of $\mathcal{O}(v)$ dictionaries, the total amount of computations is $\mathcal{O}(\sum_{k=1}^{v}(knd^{2}))=\mathcal{O}(nd^{2}v^{2})$. It’s important to note that one only needs to compute those dictionaries whose co-basis are not duplicated, which costs $\mathcal{O}(ndv)$ in computing dictionaries and determining if they pass the lex-min test. Consequently, the complexity of the method from Moss [25] in VE is $\mathcal{O}(ndv+nd^{2}v^{2}+ndv)=\mathcal{O}(nd^{2}v^{2})$.
As for the complexity of our algorithm, at each entry of a dictionary, testing a valid reverse of the Zero rule requires at most $\frac{1}{2}(n-d+n)d$ examination, which is in $\mathcal{O}(nd)$. $\mathcal{O}(v-v_{d})$ dictionaries need to test the validity, while $\mathcal{O}(v_{d})$ do not need. Therefore, the complexity is approximately given by
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d}).$ (8)
The proportion of vertex $v_{d}$ in the overall vertex has a significant impact on the actual complexity. Below, we estimate $v_{d}$ by considering two cases: one is a special case, and the other is a simple arrangement. Before estimation, we first characterize what kinds of dictionaries lie in the $d$-th layer.
Proposition 5 (Disjoint layer d).
Assume that the termination dictionary about the Zero rule is with the co-basis $N_{0}$, then those dictionaries with $N\cap N_{0}=\emptyset$ are in the $d$-th layer.
Proof: If those dictionaries with $N\cap N_{0}=\emptyset$ are not in the $d$-th layer, after repeating applying the Zero rule on the dictionary to termination, there remains at least one index in $N$ such that it is also in the co-basis of termination, which is $N_{0}$. This causes a contradiction.
$\Box$
Figure 8: LHS: a sequence of (hyper)cubes. RHS: a simple arrangement.
In the case of parallel hyperplanes, where hyperplanes form boundaries of hypercubes arranged along a specific coordinate axis, such as $y_{1},y_{2},\ldots,y_{d}=0,1,y_{1}=2,y_{1}=3,y_{1}=4,\ldots,y_{1}=n-2d+1$. As shown in the LHS of Figure 8, $v_{d}=\frac{v}{2^{d}}$, then Eq. (8) can be simplified to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})=\mathcal{O}((1-\frac{1}{2^{d}})n^{2}d^{2}v+\frac{1}{2^{d}}ndv)=\mathcal{O}((1-\frac{1}{2^{d}})n^{2}d^{2}v).$
If one stacks multiple layers of hypercubes on top of each other along this line, the proportion of $v_{d}$ will increase. Consequently, as the number of hypercubes increases, the complexity will significantly decrease.
In the case of a simple arrangement, in which every $d$ hyperplane in the arrangement form a vertex and every vertex is non-degenerate (see the RHS of Figure 8, $v=\binom{n}{d},v_{d}\geq\binom{n-d}{d}$), then Eq.(8) can be simplified to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$ $=\mathcal{O}\left(n^{2}d^{2}\left(\binom{n}{d}-\binom{n-d}{d}\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{n^{d}-(n-d)^{d}}{d!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{d^{2}n^{d-1}}{d!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(d\frac{n^{d-1}}{(d-1)!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{3}\binom{n}{d-1}+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(nd^{4}\binom{n}{d}+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(nd^{4}\binom{n}{d}\right)=\mathcal{O}(nd^{4}v).$
In a specific arrangement, because the proportion of $v_{d}$ is different, the overall computational complexity needs to be analyzed separately. But the upper bound of the computational complexity is $\mathcal{O}(n^{2}d^{2}v)$, which is significantly smaller than $\mathcal{O}(nd^{2}v^{2})$ in Moss [25]. In brief, the algorithm developed by Avis and Fukuda [4] cannot guarantee enumerating all vertices, and the algorithm by Moss [25] costs $\mathcal{O}(nd^{2}v^{2})$, while our algorithm cost at most $\mathcal{O}(n^{2}d^{2}v)$. We reasonably contend that our algorithm is state-of-the-art in the problem of VE of an arrangement.
The lower bound of the computational complexity of the VE algorithm is $\mathcal{O}(ndv)$, since at least given each vertex, one needs to take $\mathcal{O}(nd)$ to obtain another dictionary. Our algorithm achieves this lower bound at $v_{d}$ vertices. This is because the valid reverse examination is no longer needed for vertices in the $d$-th layer. The Zero rule can guarantee the termination in $d$ steps. Can we further reduce the computational complexity of the proposed algorithm? It is difficult. To do so, one may seek a new pivot rule that either increases the portion of dictionaries that waive a valid reverse examination or reduces the number of examinations per valid reverse test. In the first situation, the potential superior rule needs to have pivot steps in $\mathcal{O}(\log(d))$ or $\mathcal{O}(1)$ to terminate. However, this requirement exceeds the capabilities of polynomial pivot rules, as there is always a certain dictionary that terminates in at least $d$ steps i.e., those dictionaries discussed in Proposition 5. In the second situation, reducing the examination means less information from a dictionary but more pivot steps. Given these challenges, the proposed Zero rule is a good choice for the vertex enumeration of an arrangement.
Figure 9: A representative example to show the superiority of the Zero rule over the Criss-Cross rule.
5 Example
In this section, we provide a simple yet representative example to demonstrate the superiority of using the Zero rule. This example exposes the weaknesses of the Criss-Cross rule: First, it contains a degenerate point that will make the VE algorithm with the Criss-Cross rule go into loops. Second, the Criss-Cross rule cannot guarantee the convergence in $d$ steps. Third, the Criss-Cross rule’s convergence is not unique, which forms multiple search trees. In contrast, the VE algorithm with the Zero rule is free of loops and can find all vertices. It also only forms a single tree with the depth of $d$. In this example, we slightly abuse the dictionary and the equation between slack variables interchangeably for convenience. It is worth noting that, even though the example we have chosen demonstrates many drawbacks of the original algorithm, it is not particularly special. The coefficients of the hyperplane and their corresponding constant terms are all randomly generated, and many of the properties reflected in this example can also be found in others.
Let us consider the arrangement in Figure 9, which has 5 hyperplanes, and the coefficients and biases of each hyperplane are
$\begin{cases}&\mathbf{c}_{1}=(1,3),~{}~{}~{}~{}~{}b_{1}=4\\ &\mathbf{c}_{2}=(5,1),~{}~{}~{}~{}~{}b_{2}=5\\ &\mathbf{c}_{3}=(3,2),~{}~{}~{}~{}~{}b_{3}=2\\ &\mathbf{c}_{4}=(-1,-3),~{}b_{4}=1\\ &\mathbf{c}_{5}=(-2,1),~{}~{}~{}b_{5}=-\frac{1}{2},\end{cases}$ (9)
and slack variables are $x_{1},...,x_{5}$:
$\begin{cases}&x_{1}=4-y_{1}-3y_{2}\\ &x_{2}=5-5y_{1}-y_{2}\\ &x_{3}=2-3y_{1}-2y_{2}\\ &x_{4}=1+y_{1}+3y_{2}\\ &x_{5}=-\frac{1}{2}+2y_{1}-y_{2}.\end{cases}$
The Zero rule: Taking $x_{1}$ and $x_{2}$ as co-basis and adding the variable $x_{g}$, we obtain the following initial dictionary:
$\begin{cases}x_{3}&=-\frac{5}{2}x_{g}+\frac{1}{2}x_{1}+\frac{1}{2}x_{2}\\ x_{4}&=5x_{g}-x_{1}+0x_{2}\\ x_{5}&=0x_{g}+\frac{1}{2}x_{1}-\frac{1}{2}x_{2}.\end{cases}$ (10)
The dictionary with $N=\{g,1,2\}$ holds the lexicographically smallest order among all co-basis in this arrangement. However, this dictionary fails the lex-min test, which means we will not use this dictionary to produce a vertex.
Next, we examine each entry and select the one with the smallest index that satisfies the condition of the valid reverse. $(3,1)$ in Eq. (10) is not a valid reverse, while $(4,1)$ is. Therefore, we pivot $(4,1)$ and obtain the dictionary below:
$\begin{cases}x_{1}&=5x_{g}+0x_{2}-x_{4}\\ x_{3}&=0x_{g}+\frac{1}{2}x_{2}-\frac{1}{2}x_{4}\\ x_{5}&=\frac{5}{2}x_{g}-\frac{1}{2}x_{2}-\frac{1}{2}x_{4}.\end{cases}$ (11)
This dictionary also fails the lex-min test. We need to continue testing reverses based on this dictionary. $(1,2)$ is invalid, and $(3,2)$ is valid. Therefore, we pivot (3,2) and obtain the following dictionary:
$\begin{cases}x_{1}&=5x_{g}+0x_{3}-x_{4}\\ x_{2}&=0x_{g}+2x_{3}+x_{4}\\ x_{5}&=\frac{5}{2}x_{g}-x_{3}-x_{4}.\end{cases}$ (12)
This dictionary successfully passes the lex-min test. Therefore, we can output the basis $B=\{1,2,5\}$. The current dictionary is at the $2$-nd layer, and there is no need to proceed further. As a result, we can skip checking the valid reverses from this dictionary and pivot (2,3) to return to Eq. (11).
After returning to Eq. (11), we only need to search entries whose row numbers or column numbers are larger than 3. While $(5,2)$, $(1,4)$, $(3,4)$, $(5,4)$ are all invalid reverses, Eq. (11) is finished. We can apply the Zero rule on it and go to the dictionary Eq. (10). Next, we continue testing and repeat this process. The LHS of Figure 10 presents a spanning tree that encompasses all dictionaries and their orders in the reverse process using the Zero rule.
Figure 10: Spanning trees that connect all vertices obtained by the VE algorithm. LHS: Zero rule. RHS: the Criss-Cross rule.
The Criss-Cross rule: Now, we utilize the Criss-Cross rule to solve this small example. Note that starting from any dictionary, repetitively applying the Criss-Cross rule will eventually lead to an optimal dictionary. Thus, the initial dictionary chosen for the VE algorithm is considered as an optimal dictionary. However, the optimal dictionary is not unique. In this example, we set the initial dictionary as Eq. (13), which is the same as the Zero rule.
$\begin{cases}x_{3}&=-\frac{5}{2}x_{g}+\frac{1}{2}x_{1}+\frac{1}{2}x_{2}\\ x_{4}&=5x_{g}-x_{1}+0x_{2}\\ x_{5}&=0x_{g}+\frac{1}{2}x_{1}-\frac{1}{2}x_{2}\\ x_{f}&=0x_{g}-x_{1}-x_{2}.\end{cases}$ (13)
Meanwhile, another optimal dictionary is
$\begin{cases}x_{1}&=0x_{g}+x_{2}+2x_{5}\\ x_{3}&=-\frac{5}{2}x_{g}+x_{2}+x_{5}\\ x_{4}&=5x_{g}-x_{2}-2x_{5}\\ x_{f}&=0x_{g}-x_{2}-2x_{5}.\end{cases}$ (14)
We need to apply the Zero rule based on these two dictionaries, respectively. Let us start from Eq. (13). Similarly, this dictionary does not pass the lex-min test, and $(4,1)$ is a valid reverse pivot about the Criss-Cross rule. Pivoting $(4,1)$ can obtain the following dictionary:
$\begin{cases}x_{1}&=5x_{g}+0x_{2}-x_{4}\\ x_{3}&=0x_{g}-\frac{1}{2}x_{2}+\frac{1}{2}x_{4}\\ x_{5}&=\frac{5}{2}x_{g}-x_{2}-\frac{1}{2}x_{4}\\ x_{f}&=-5x_{g}-0x_{2}+2x_{4}.\end{cases}$ (15)
which does not pass the lex-min test, either.
In Eq. (15), $(5,2)$ is a valid reverse pivot according to the Criss-Cross rule. Pivoting it results in the dictionary shown in Eq. (16). This new dictionary successfully passes the lex-min test, and we can compute the vertex from it.
$\begin{cases}x_{1}&=5x_{g}-x_{4}+0x_{5}\\ x_{2}&=5x_{g}-x_{4}-2x_{5}\\ x_{3}&=-\frac{5}{2}x_{g}+x_{4}+x_{5}\\ x_{f}&=-10x_{g}+2x_{4}+2x_{5}.\end{cases}$ (16)
Upon examining Eq. (16), we find that there are no valid reverses according to the Criss-Cross rule. Next, we apply the Criss-Cross rule to this dictionary, which identifies $(3,4)$. However, this step initiates an endless cycle in the VE algorithm, involving dictionaries with co-basis $\{4,5\}$, $\{3,5\}$, $\{1,3\}$, $\{1,2\}$, and $\{1,5\}$. The cycle is depicted in the RHS of Figure 10.
Next, let us apply the Criss-Cross rule on Eq. (14). No entries are valid reverses; therefore, Eq. (14) is an isolated dictionary.
6 Systematic Experiments
Here, we conduct systematic experiments to demonstrate the superiority of the proposed algorithm using the Zero rule over algorithms in Avis and Fukuda [4] and Moss [25]. We independently implement the algorithms in Avis and Fukuda [4] and Moss [25] by following their algorithmic structures. We implement our algorithm according to the outline shown in Algorithm 6. We generate four different arrangements: unit hypercubes, unit cubes with a cone being cut i.e., adding the equation $y_{1}+y_{2}+\cdots+y_{d}=1.5$ with the unit cube, random hyperplanes defined by random integers as coefficients, and a three-layer ReLU neural network. The merit of the former is the ground truth is known, which can assist us to examine if the compared algorithms can enumerate all vertices upon completion of the algorithm. The latter is to show the comparison in the general case. All our examples and Matlab code are publicly available [^1] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm for readers’ free download and use.
Unit hypercubes: Table 2 presents different algorithms’ performance over unit hypercubes ranging from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$. It is surprising that Avis and Fukuda [4]’s algorithm cannot find all vertices in such a direct case. The reason is that this algorithm skips many vertices and stops early. Specifically, in certain steps, applying the Criss-Cross rule jumps into the initial dictionary and selects much later entries to pivot, which makes the algorithm terminate earlier without finding all vertices.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
Truth
16
32
64
128
256
512
1024
Moss [25]
16
32
64
128
256
512
1024
Avis and Fukuda [4]
5
6
7
8
9
10
11
Ours
16
32
64
128
256
512
1024
Table 2: The number of vertices found by different algorithms over unit hypercubes from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$.
Unit hypercubes with a cone being cut: Table 3 presents different algorithms’ performance over unit hypercubes with a cone being cut ranging from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$. In these examples, Avis and Fukuda’s algorithm [4] once again fails to find all vertices. Unlike the previous performance, their algorithm goes into a loop, causing it to fail in finding all vertices. After the table, Figures 12 and Figure 12 demonstrate the process of different algorithms in finding all vertices in $\mathbb{R}^{9}$ and $\mathbb{R}^{10}$, respectively; our method is better.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
Moss [25]
48
112
256
576
1280
2816
6144
Avis and Fukuda [4]
6
7
8
9
10
11
12
Ours
48
112
256
576
1280
2816
6144
Table 3: The number of vertices found by different algorithms over unit hypercubes with a cone being cut from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$.
Figure 11: cube with a cone being cut in $\mathbb{R}^{9}$
Figure 11: cube with a cone being cut in $\mathbb{R}^{9}$
Figure 12: cube with a cone being cut in $\mathbb{R}^{10}$
Random arrangement. By taking random integers as coefficients of arrangements, a total of 25 cases are generated by varying from $\mathbb{R}^{2}$ to $\mathbb{R}^{6}$ and from 7 to 15 hyperplanes. Each case has five runs to ensure the reliability of the results. Table 4 displays the average number of vertices found by three algorithms. Table 5 shows the average computation time. For each case, we set 5000 seconds as the critical value. If the time in a case exceeds this value, we denote it as $>5000$. The algorithm falls into a cycle in processing all examples when using Avis and Fukuda’s method.
$d\backslash n$
Method
7
9
11
13
15
$\mathbb{R}^{2}$
Moss [25]
21
35
53.8
75.2
99.6
Avis and Fukuda [4]
8.2
7.4
8.6
7
8
Ours
21
35
53.8
75.2
99.6
$\mathbb{R}^{3}$
Moss [25]
35
83.8
163
283.8
445.2
Avis and Fukuda [4]
9
8.2
8.8
16.4
11.2
Ours
35
83.8
163
283.8
445.2
$\mathbb{R}^{4}$
Moss [25]
35
126
330
714
1364.4
Avis and Fukuda [4]
9
9.8
12.4
12
11
Ours
35
126
330
714
1364.4
$\mathbb{R}^{5}$
Moss [25]
21
126
462
1287
3002.6
Avis and Fukuda [4]
7.6
8.6
19.2
10.4
14
Ours
21
126
462
1287
3002.6
$\mathbb{R}^{6}$
Moss [25]
7
84
462
1716
5005
Avis and Fukuda [4]
4.6
10.4
10.2
8.6
13.6
Ours
7
84
462
1716
5005
Table 4: The average number of vertices found by different algorithms.
$d\backslash n$
Method
7
9
11
13
15
$\mathbb{R}^{2}$
Moss [25]
1.8357
3.4629
6.3672
10.339
15.709
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
1.7081
3.5555
6.5597
10.531
16.288
$\mathbb{R}^{3}$
Moss [25]
2.8036
8.2725
19.666
39.980
61.593
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
2.6670
8.0951
19.207
39.196
70.270
$\mathbb{R}^{4}$
Moss [25]
2.6872
12.382
40.707
102.05
229.00
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
22.5885
12.051
38.789
100.72
217.85
$\mathbb{R}^{5}$
Moss [25]
1.4985
12.137
55.207
186.65
502.96
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
1.4672
12.059
54.835
184.75
483.07
$\mathbb{R}^{6}$
Moss [25]
0.4936
7.6735
54.772
246.42
960.63
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
0.4702
7.9303
54.968
246.02
817.43
Table 5: The average computation time consumed by different algorithms.
Figure 13: average number of vertex $\mathbb{R}^{5}$
Figure 13: average number of vertex $\mathbb{R}^{5}$
Figure 14: average number of vertex $\mathbb{R}^{6}$
Using the Zero rule significantly improves Moss’ method, while reversing the Criss-Cross rule did not finish all cases well. Additionally, from each column in Table 5, Figure 14 and Figure 14, it can be observed that as the number of dimensions and vertices increases, the improvement becomes even more pronounced for any number of hyperplanes. This aligns with our complexity analysis, indicating that if the number of hyperplanes and the number of dimensions continuously increase, the Zero rule can save more computational overhead.
A three-layer ReLU neural network: We have built a neural network model with an input layer consisting of $7$ nodes to handle data from $\mathbb{R}^{5}$. Each node in the input layer will receive a $5$-dimensional input. The hidden layer comprises $5$ nodes, each utilizing a weight vector with $7$ dimensions along with a bias term. These layers employ the Rectified Linear Unit (ReLU) activation function to introduce non-linearity into the network. In this network, all weight vectors and bias terms are randomly generated using the rand function in MATLAB. Firstly, we use auxiliary linear programming to find the partition of the seven neurons in the input layer, resulting in a total of $119$ cells in $\mathbb{R}^{5}$. As the activation state of each cell is different, $5$ hyperplanes corresponding to the nodes in the hidden layer will also be different. Next, we use algorithms to find out the total number of vertex of these $119$ different hyperplane arrangements with $12=7+5$ hyperplanes in $\mathbb{R}^{5}$, meanwhile, during the execution of each algorithm, we also select and count the vertices on the cell that corresponding to the activation states and the time taken for these computations. The naming order of the cells is only related to the sequence in which we found them, and partly of the results are provided in Tables 6 and 7 (Moss: Moss [25]; AF: Avis and Fukuda [4]) and the overall result can be found in our link [^2] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm, $-1$ in the first cell means none of the seven nodes in the input layer is active.
Method 1 4 7
feasible
all
feasible
all
feasible
all
0
[25]
$-1$
$-1$
7
246
4
246
[4]
$-1$
$-1$
1
12
1
17
Ours
$-1$
$-1$
7
246
4
246
10
[25]
9
246
2
456
7
96
[4]
1
7
1
8
1
7
Ours
9
246
2
456
7
96
20
[25]
5
246
6
666
10
456
[4]
1
9
1
16
1
6
Ours
5
246
6
666
10
456
30
[25]
8
666
3
246
1
666
[4]
1
11
2
16
1
7
Ours
8
666
3
246
1
666
40
[25]
6
666
3
666
4
456
AF
2
3
2
11
2
8
Ours
6
666
3
666
4
456
50
[25]
2
792
5
666
4
792
[4]
1
17
2
6
1
13
Ours
2
792
5
666
4
792
60
[25]
6
246
3
246
4
666
[4]
3
14
1
7
2
8
Ours
6
246
3
246
4
666
70
[25]
9
666
4
792
9
456
[4]
3
18
2
7
5
21
Ours
9
666
4
792
9
456
80
[25]
7
666
10
666
7
666
[4]
3
8
2
6
2
7
Ours
7
666
10
666
7
666
90
[25]
4
246
5
666
2
792
[4]
1
11
2
3
1
11
Ours
4
246
5
666
2
792
100
[25]
7
792
3
792
4
666
[4]
1
6
1
6
2
4
Ours
7
792
3
792
4
666
110
[25]
3
792
8
792
5
792
[4]
2
10
2
7
1
22
Ours
3
792
8
792
5
792
Table 6: The number of feasible and all vertices in corresponding arrangement found by different algorithms.
Method
1
4
7
11
14
17
0
Moss
$-1$
39.224
36.991
36.848
69.985
13.7739
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
$-1$
34.216
37.636
35.147
67.154
14.3105
20
Moss
36.215
103.53
70.071
102.83
36.260
102.93
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
36.252
93.205
66.960
95.646
35.053
94.164
40
Moss
104.52
103.48
69.086
127.26
104.40
127.10
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
93.487
93.798
67.129
110.63
94.500
112.92
60
Moss
36.556
38.099
104.10
105.85
125.58
70.180
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
37.892
36.202
94.394
96.062
111.79
67.700
80
Moss
104.75
104.90
104.78
36.840
104.45
125.81
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
95.837
97.517
96.017
36.358
93.960
111.37
100
Moss
125.51
125.95
104.36
126.42
127.87
126.05
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
110.67
110.24
95.478
110.19
111.65
109.63
Table 7: The computation time consumed by different algorithms.
From Table 6, it can be seen that the vertices in these hyperplane arrangements are highly degenerate. In a simple arrangement formed by 12 hyperplanes in $\mathbb{R}^{5}$, there should be $492$ vertices. However, only a small portion of these examples reach this value. Therefore, we can also observe the effectiveness of our method in highly degenerate hyperplane distributions from this series of examples. The following Figures 16-22 demonstrate the process of searching vertices in four randomly selected examples, those figures show that when searching vertices of arrangement, our algorithm is better than the random generate integer coefficient arrangement in previous, but when searching vertices on the cell, Their performance is similar and less efficient, but Avis and Fukuda’s method, while not finding all the vertices, is highly efficient in generating vertices on the cell.
Figure 15: vertices in 4-th arrangement
Figure 15: vertices in 4-th arrangement
Figure 16: vertices in 4-th ell
Figure 17: vertices in 47-th arrangement
Figure 18: vertices in 47-st arrangement
Figure 19: vertices in 81-st arrangement
Figure 20: number of vertex in 81-st cell
Figure 21: vertices in 114-th arrangement
Figure 17: vertices in 47-th arrangement
Figure 18: vertices in 47-st arrangement
Figure 19: vertices in 81-st arrangement
Figure 20: number of vertex in 81-st cell
Figure 21: vertices in 114-th arrangement
Figure 22: vertices in 114-th cell
7 Discussion
The above experiments demonstrate that the VE algorithm using the Criss-Cross rule cannot enumerate all vertices. Can we remedy this problem? The reason why using the Criss-Cross rule fails in certain cases is the unrigorous examination of valid reversibility. Here, we complement the Criss-Cross rule by upgrading its if condition to the if-and-only-if condition.
Proposition 7.1.
Let $x_{B}=\bar{A}x_{N}$ be an arbitrary dictionary with the objective function. Then $(s,r)$, where $s\in B_{\neq f}$ and $r\in N_{\neq g}$, is a valid reverse pivot of the Criss-Cross rule if and only if either (a)
$(a1)~{}\bar{a}_{sg}>0,\bar{a}_{sr}>0,\bar{a}_{sj}\geq 0\ for\ j\in N_{\neq g},j<s.$
$(a2)~{}\forall j<r,\ \mathrm{If}\ j\in B_{\neq f},\bar{a}_{jg}\bar{a}_{sr}\geq\bar{a}_{jr}\bar{a}_{sg};\ \mathrm{If}\ j\in N_{\neq g},\bar{a}_{fr}\bar{a}_{sj}\geq\bar{a}_{fj}\bar{a}_{sr}.$
$(a3)~{}\textup{ If }s<r,\bar{a}_{fr}\leq 0.$
or (b)
$(b1)~{}\bar{a}_{fr}<0,\bar{a}_{sr}<0,\bar{a}_{ir}\leq 0\ \mathrm{for}\ i\in B_{\neq f},i<r$
$(b2)~{}\forall i<s,\ \mathrm{If}\ i\in B_{\neq f},\bar{a}_{ir}\bar{a}_{sg}\geq\bar{a}_{ig}\bar{a}_{sr};\ \mathrm{If}\ i\in N_{\neq g},\bar{a}_{fi}\bar{a}_{sr}\geq\bar{a}_{fr}\bar{a}_{si}.$
$(b3)~{}\mathrm{If}~{}r<s,\bar{a}_{sg}\geq 0.$
holds.
Proof: Let $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ be the dictionary resultant from pivoting $(s,r)$ on $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$.
$\Rightarrow$: Assume $(s,r)$ be a valid reverse Criss-Cross pivot. Then the Criss-Cross rule pinpoint $(r,s)$ on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ to pivot. Assume $i$ to be the smallest index such that $x_{i}$ is dual or primal infeasible, we study cases $i\in\tilde{B}$ and $i\in\tilde{N}$ separately.
Case 1. Let $i\in\tilde{B}$, then $r=i$ and $\tilde{a}_{rg}<0$, $s$ be the smallest index such that $\tilde{a}_{rs}>0$.
(a1): $\bar{a}_{sr}=\frac{1}{\tilde{a}_{rs}}>0$ and $\bar{a}_{sg}=-\tilde{a}_{rg}\bar{a}_{rs}>0$. Since $\forall j\in\tilde{N}_{\neq g}$ and $j<s$, $\tilde{a}_{rj}\geq 0$, then $\bar{a}_{sj}=-\tilde{a}_{rj}\bar{a}_{rs}\geq 0$. Note that $\bar{a}_{sr}\geq 0$, then $\forall j\in N_{\neq g}$ and $j<s,\bar{a}_{sj}\geq 0$, (a1) holds.
(a2): $\forall j<r,x_{j}$ is either primal or dual feasible in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$.
Note that as $j=s$, $\bar{a}_{jg}\bar{a}_{sr}=\bar{a}_{jr}\bar{a}_{sg}$. Also note that for those $j\in\tilde{B}_{\neq f}$ and $j<r$, $x_{j}$ is primal feasible on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, then $\tilde{a}_{jg}=\bar{a}_{jg}-\frac{\bar{a}_{jr}\bar{a}_{sg}}{\bar{a}_{sr}}\geq 0$ means $\bar{a}_{jg}\bar{a}_{sr}\geq\bar{a}_{jr}\bar{a}_{sg}$ $\forall j\in B_{\neq f}$ and $j<r$. Since $\forall j\in\tilde{N}_{\neq g}$ and $j<r$, $x_{j}$ dual feasible on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, then $\tilde{a}_{fj}=\bar{a}_{fj}-\frac{\bar{a}_{fr}\bar{a}_{sj}}{\bar{a}_{sr}}\leq 0$ means $\bar{a}_{fr}\bar{a}_{sj}\geq\bar{a}_{fj}\bar{a}_{sr}$. (a2) holds.
(a3): (a1-a2) are sufficient for $s>r$. In the case of $s<r$, $x_{s}$ on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ is dual feasible, $\tilde{a}_{fs}\leq 0$, and $\bar{a}_{fr}=\tilde{a}_{fs}\bar{a}_{sr}\leq 0$. (a3) holds.
Case 2. Let $i\in\tilde{N}$, then $s=i$ and $\tilde{a}_{fs}>0$, $r$ be the smallest index such that $\tilde{a}_{rs}<0$. Similar to the study of the first case, (b1-b3) can be obtained.
$\Leftarrow$: Assume one of the (a) and (b) holds.
Case 1. Let the part (a) hold. Note that $\forall j\in\tilde{B}_{\neq f}$ and $j<r$, by (a1) and (a2), $\tilde{a}_{jg}=\bar{a}_{jg}-\frac{\bar{a}_{jr}\bar{a}_{sg}}{\bar{a}_{sr}}=\frac{1}{\bar{a}_{sr}}(\bar{a}_{jg}\bar{a}_{sr}-\bar{a}_{jr}\bar{a}_{sg})\geq 0,x_{j}$ on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ is primal feasible. Also $\forall j\in\tilde{N}_{\neq g}$ and $j<r$, if $s\geq r$, by (a2), $\tilde{a}_{fj}=\bar{a}_{fj}-\frac{\bar{a}_{fr}\bar{a}_{sj}}{\bar{a}_{sr}}=-\frac{1}{\bar{a}_{sr}}(\bar{a}_{fr}\bar{a}_{sj}-\bar{a}_{fj}\bar{a}_{sr})\leq 0,x_{j}$ is dual feasible on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$. If $s<r$, for the additional case $j=s$, by (a3), $\tilde{a}_{fs}=\frac{\bar{a}_{fr}}{\bar{a}_{sr}}\leq 0,x_{s}$ is dual feasible on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$. Hence, for any $j<r$, $x_{j}$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ is dual or primal feasible.
Since $\tilde{a}_{rg}=-\frac{\bar{a}_{sg}}{\bar{a}_{sr}}<0$, $x_{r}$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\tilde{N}$ is primal infeasible, the Criss-Cross rule pinpoint at $r\in\tilde{B}_{\neq f}$. Since $\forall j\in\tilde{N}_{\neq g}$ and $j<s$, by (a1), $\tilde{a}_{rj}=-\frac{\bar{a}_{sj}}{\bar{a}_{sr}}\leq 0$ and $\tilde{a}_{rs}=\frac{1}{\bar{a}_{sr}}>0$, thus Criss-Cross rule gives $s\in\tilde{N}$.
Hence, the Criss-Cross rule is fixed $(r,s)$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, and $(s,r)$ is a valid reverse Criss-Cross rule pivot.
Case 2. Let the part (b) hold. Similar to the first case, applying the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ the first step fix at $s\in\tilde{N}_{\neq g}$ and second step pinpoint $r$ in the column of $\mathbf{x}_{s}$, it follows $(s,r)$ is still valid reverse Criss-Cross rule pivot.
$\Box$
Table 8 shows that the complemented Criss-Cross rule can find all vertices using the same hypercubes in Table 2.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
Truth
16
32
64
128
256
512
1024
After Complement
16
32
64
128
256
512
1024
Original
5
6
7
8
9
10
11
Table 8: The result from complete and incomplete reverse Criss-Cross rule.
Compared with the if condition in [4], our complement exerts two more restrictions. But the complexity of examining the valid reverse of the Criss-Cross rule is still $\mathcal{O}(n)$. Therefore, the total complexity is $\mathcal{O}(n^{2}dv)$. In contrast, the complexity of the Zero rule can be as low as $\mathcal{O}(nd^{4}v)$, which means that the Zero rule is suitable for an arrangement with more hyperplanes, while the Criss-Cross rule is appropriate for an arrangement with higher dimensions.
8 Conclusion
In this paper, we present the Zero rule for the VE problem of hyperplane arrangements. Compared to the classical Criss-Cross rule, the Zero rule enjoys several computationally friendly properties: i) It can get rid of objective functions, which saves the computation. ii) The condition between the Zero rule and its valid reverse is sufficient and necessary, which prevents the algorithm from going into a loop and catastrophically skipping legitimate dictionaries. iii) In $\mathbb{R}^{d}$, its pivot steps are at most $d$, which can transform the unpredictable depth into a predictable maximum depth of $d$. iv) Its termination dictionary within the entire arrangement is unique. This approach, originating from the arrangement of hyperplanes itself, also contributes to establishing rules with fewer convergence points on a polytope. Because of these properties, the complexity of the VE algorithm using the Zero can be as low as $\mathcal{O}(nd^{4}v)$. Systematic and comprehensive experiments have confirmed that the Zero rule is the state of the art in the VE problem of hyperplane arrangements. In the future, more efforts will be made to increase the proportion of optimal solutions in the dictionary or reduce the amount of information needed to verify reversibility to further escalate the algorithmic efficiency.
9 Acknowledgement
The authors are grateful for Prof. Ning Cai’s suggestions (Hong Kong University of Science and Technology at Guangzhou), discussions with Mr. Xiangru Zhong, a visiting graduate student at the Department of Mathematics, The Chinese University of Hong Kong, and anonymous reviewers’ advice.